OpenStudy (anonymous):

Okay, another integral. (problem inside) Hints. don't do my work:) I'll see what I can do and what I can't.

OpenStudy (anonymous):

\[\Large \int\limits_{ }^{ } \frac{1}{\sqrt{\theta-1}+\sqrt{(\theta-1)^3}}d\theta\]

OpenStudy (anonymous):

Well, firstly, I would substitute, \[\LARGE w=\theta-1~~~~~dw=d \theta\](tired of U s)

OpenStudy (anonymous):

\[\Large \int\limits_{ }^{ } \frac{1}{\sqrt{w}+\sqrt{w^3}}dw\]

OpenStudy (darkprince14):

factor, and use substitutions. check the pattern.

OpenStudy (anonymous):

Factor?

OpenStudy (anonymous):

Can we come back to the factoring later, ones I finish, if I be able to?

OpenStudy (darkprince14):

the denominator, that was just a random guess...II haven't even started solving it myself :)

OpenStudy (anonymous):

and then what, partial fraction?

OpenStudy (kainui):

I would probably pick sqrt(theta-1)=u instead.

OpenStudy (anonymous):

Ohh, okay, I'll try this.

OpenStudy (anonymous):

\[\Large \int\limits_{ }^{ } \frac{1}{(\sqrt{\theta -1 }+\sqrt{\theta -1 }(\theta -1 )^2}d\theta\] but there will be a difficulty with du/d(theta) part though

OpenStudy (anonymous):

Oops, the ^2 shouldn't be there.

OpenStudy (anonymous):

Let's first do the w sub. we can always come back if we want, can't we?

OpenStudy (anonymous):

\[\Large \int\limits\limits_{ }^{ } \frac{1}{(\sqrt{w}+w\sqrt{w})}dw\]

OpenStudy (anonymous):

it is just looks better.

OpenStudy (kainui):

Well there's more than one way to do this problem, I think after the substitution I've done it becomes much nicer and obvious.

OpenStudy (anonymous):

I'll try partial fractions. I am new to it, but I saw a couple examples from the book. We didn't learn this method in class yet.

OpenStudy (kainui):

I would probably rewrite your square roots as w^1/2 instead. I did it without partial fractions.

OpenStudy (anonymous):

\[\Large \int\limits\limits_{ }^{ } \frac{A}{(\sqrt{w}}+\frac{B}{w}dw\]\[\Large A(w)+B(\sqrt{w})=1\] \[\Large A+B=1\]\[\Large 4A+2B=1\]

OpenStudy (anonymous):

I plugged x=1, and then x=4.

OpenStudy (anonymous):

\[\Large B=1-A\]

OpenStudy (kainui):

I don't think you can do partial fractions like that.

OpenStudy (anonymous):

yes I can, I am fairly, sure that the A(w)+B(sqrt w) =1 is true for all values.

OpenStudy (anonymous):

MY dad is calling, I got to go. Sorry for disappointing y'all.

OpenStudy (anonymous):

I'll close it for now.

OpenStudy (kainui):

I'll write up my answer, and you can see what I did later.

OpenStudy (zarkon):

\[\Large \int\frac{1}{(\sqrt{w}+w\sqrt{w})}dw=\Large \int\frac{1}{\sqrt{w}}\frac{1}{(1+(\sqrt{w})^2)}dw\]

OpenStudy (zarkon):

\[=\Large \int\frac{1}{2\sqrt{w}}\frac{2}{(1+(\sqrt{w})^2)}dw\]

OpenStudy (kainui):

\[\Large u = (\theta -1)^\frac{1}{2} \\ \Large 2du = \frac{1}{(\theta - 1 )^\frac{1}{2}}d \theta\] now factor the integral like this: \[\Large \int\limits \frac{1}{1+(\theta-1)} \frac{d \theta}{ (\theta-1)^\frac{1}{2}}\] Now be careful plugging it in and pay attention to what I've done: \[\Large \int\limits \frac{1}{1+u^2} 2 du\] You should recognize this integral. =P